# Cluster analysis of earthquake hypocenters in Azerbaijan and surrounding territories

### v1

@@ 6 Analysis of the earthquake depth distributions
Thus, let us construct the earthquake depth histograms for the clusters shown in Fig.4b. For comparison, the histogram of all earthquakes is also depicted in Fig.7. All histograms exhibit a maximum around 10 km, with the upper cutoff depth at 60 km and lower cutoff depths for the clusters. This is in good agreement with other seismic research (e.g., [6]) and the assessment of Moho depth in Azerbaijan, which, according to [4], varies from approximately 40 km in the Kura basin and northern foreland basin to over 50 km in the eastern Greater Caucasus.
As shown by [44], to fit the histograms (see details in Appendix B), the probability density functions (PDFs) of a mixture of normal distributions are used. The application of this PDF fitting procedure revealed that attempts to approximate the histograms in Figs.7c and 7d using a mixture of normal distributions were unsuccessful. Instead, a single normal distribution was found to be more appropriate. Interestingly, the histograms in Figs.7e and 7f exhibit a bimodal nature. Numerical verification shows that the PDF for the histogram of the total number of earthquake depths has two modes, located at approximately 10 km (intense) and 30 km (almost indistinguishable). The latter peak becomes more pronounced in the cluster histograms (Figs.7e, f). In particular, the deeper peak in cluster 6 (Fig.7f) is located at a depth of around 35 km. The studies by [50] on seismicity in the Ismayilly area, which is covered by cluster 6, also confirm the accumulation of hypocenters at a depth of about 40 km.
As indicated by [49], the two modes observed in the depth histogram for Japanese earthquakes reflect distinct sources: the shallow peak (around 5 km) is associated with crustal earthquakes and anthropogenic events, while the deeper peak (around 35 km) is attributed to seismic events near the plate boundary and deep crustal earthquakes. Other factors contributing to the formation of these two modes may be related to the presence of complex tectonic structures, as discussed by [2] and [50]. Overall, this issue requires a comprehensive analysis of geological and geophysical data for the region.
@@ 7 Concluding remarks
Using the DBSCAN and HDBSCAN algorithms, we evaluated partitions of the earthquake catalog. Since there are no strict restrictions on the intrinsic parameters of the algorithms, we obtained a variety of partitions containing between 2 and 20 clusters. To assess the quality of clustering and select the appropriate partitions for further study, the Silhouette Index was calculated. Partitions providing the maximum Silhouette Index, or those close to it, were selected. The next step in reducing the number of partitions involved using the Adjusted Rand Index. This index was calculated for pairs of partitions with the same number of clusters, allowing us to identify similar partitions. As a result of these clustering stages, we selected partitions containing 5-8 clusters, which are the most informative from a data perspective.
@@
Data availability
The earthquake catalog used in the current study is available at http://www.isc.ac.uk/iscbulletin/search/. The fault catalog is available at http://neotec.ginras.ru/index/english/database_eng.html. Python source code implementing the cluster analysis of the earthquake catalog is available at https://github.com/SkurativskaKateryna/IJ_NLM_Earthquake_clusterization.git.
@@ Appendix A Estimation of DBSCAN parameters
The selection of the parameters $\varepsilon$ and $Min\_samples$ for the DBSCAN algorithm is based on the analysis of the sorted $k$-distance graph, as discussed by [26] and [15], which uses the sorted distance matrix. We plot the $k$-distance graphs for $k=4$, $k=70$, and $k=150$ (Fig. 8). The appropriate value of $\varepsilon$ corresponds to the formation of the ”knee” on these graphs. Therefore, we choose $\varepsilon$ from the range $[0.2;0.6]$ and $Min\_samples$ from the range $[25;150]$. It is important to note that this procedure should still be considered an approximate method for parameter estimation, as noted by [16] and [51].
@@
Appendix B Fitting the earthquake depth histograms by a mixture of normal distributions
The histograms shown in Fig.7 are approximated by a mixture of normal distributions. To do this, we used the Mathematica 12.0 in-built procedure EstimatedDistribution[data, MixtureDistribution[{a, b}, NormalDistribution[$\mu_{1}$, $\sigma_{1}$], NormalDistribution[$\mu_{2}$, $\sigma_{2}$]]], where $data$ stands for the array of depths, while the parameters $a$, $b$ (it is obviously that $a+b=1$, but we follow the Mathematica syntax), $\mu_{1,2}$, and $\sigma_{1,2}$ should be estimated. Then the resulting PDF coincides with a linear combination of the normal distributions $N(\mu_{1,2},\sigma_{1,2})$. The calculated values of parameters are presented in Table 2.
Table 2: The parameters for the mixture of normal distributions depicted in Fig.7.
$a$ $b$ $\mu_{1}$ $\sigma_{1}$ $\mu_{2}$ $\sigma_{2}$
Total depths 0.674 0.326 9.945 4.842 30.147 15.137
Fig.7(a) 0.644 0.356 11.148 5.011 24.235 10.126
Fig.7(b) 0.937 0.063 8.140 3.749 21.719 9.272
Fig.7(c) 9.892 4.291
Fig.7(d) 9.614 4.986
Fig.7(e) 0.698 0.302 2.544 4.362 8.444 3.223
Fig.7(f) 0.732 0.268 9.166 4.849 34.938 10.027
It turned out that this procedure can fail when $b$ is close to zero. Then we applied EstimatedDistribution[data, NormalDistribution[$\mu_{1}$, $\sigma_{1}$].

### v2

@@ 6 Analysis of the earthquake depth distributions
Thus, using the partitions “db 192” and “hdb 15”, we fit a probability distribution function (PDF) to the histogram representing the depth distribution for each cluster. To do this, we choose PDF in the form of a mixture of two normal distributions, as recommended by [47], and evaluate the mixture’s parameters using the proper Python class, mentioned in Appendix B. For comparison, the histogram and corresponding PDF for total earthquake counts are also depicted in Fig.8. The PDF exhibits two maxima around 10 km and 32 km, with the upper cutoff depth at 60 km. This is in good agreement with other seismic research (e.g., [6]) and the assessment of Moho depth in Azerbaijan, which, according to [4], varies from approximately 40 km in the Kura basin and northern foreland basin to over 50 km in the eastern Greater Caucasus.
The application of the fitting procedure to each cluster of the partition “db 192” provides the parameters for the PDF curves, presented in Table 2. The parameters of PDF functions corresponding to the partition “hdb 15” are included in Table 3. The resulting PDF curves are depicted in Figs.6a and 6b. In general, the shapes of PDF profiles are mostly similar for the partitions. In Fig.6a, we can distinguish Clusters number 1, which is characterized by the bi-modal PDF. The rest of the clusters are unimodal. The cluster’s PDF maxima are observed at depth 15.31 km and 27.02; 14.72 km; 8.32 km; 6.15 km; 10.28 km; 9.11 km. The PDF profiles in Fig.6b exhibit almost similar properties. The distributions associated with clusters 3, 4, and 6 are characterized by bimodality. The PDF modes are 15.02 km and 25.1 km; 10.2956 km and 15.668 km; 10.19 km; 6.93 km; 9.74 km; 5.32 km and 10.66 km. Thus, we see that both algorithms produce four clusters (1,2,3, and 4 in Fig.6), which are nearly identical. This indicates a high degree of reliability in the classification of earthquakes in these areas. Clusters 5 and 6 in Fig.6a and the same pair in Fig.6b also contain a large number of common points. This means that the algorithms correctly identified seismic events into the clusters. The reasons for the differences in the details of the geometric shape of the clusters may be related to the peculiarities of the algorithms or to the hidden properties of the distribution of earthquakes. The latter circumstance is undoubtedly a particularly intriguing issue for further research.
A brief note should be made regarding bi-modal PDFs. As indicated in [52], the two modes observed in the depth histogram for Japanese earthquakes reflect distinct sources: the shallow peak (around 5 km) is associated with crustal earthquakes and anthropogenic events, while the deeper peak (around 35 km) is attributed to seismic events near the plate boundary and deep crustal earthquakes. Other factors contributing to the formation of these two modes may be related to the presence of complex tectonic structures, as discussed in [2] and [53]. Overall, this issue requires a comprehensive analysis of geological and geophysical data for the region.
@@ 7 Concluding remarks
Using the DBSCAN and HDBSCAN algorithms, we evaluated partitions of the earthquake catalog. Since there are no strict restrictions on the intrinsic parameters of the algorithms, we obtained a variety of partitions containing between 2 and 10 clusters. To assess the quality of clustering and select the appropriate partitions for further study, the Silhouette Index was calculated. Partitions providing the maximum Silhouette Index, or those close to it, were selected. The next step in reducing the number of partitions involved using the Adjusted Rand Index. This index was calculated for pairs of partitions with the same number of clusters, allowing us to identify similar partitions. As a result of these clustering stages, we selected partitions containing 4 or 6 clusters, which are the most informative from a data perspective.
@@
Funding The work is supported in part by the NATO Project G5907 and Project 0121U107662, National Academy of Sciences of Ukraine, Projects 0123U100183.
Data availability The earthquake catalog used in the current study is available at http://www.isc.ac.uk/iscbulletin/search/. The fault catalog is available at http://neotec.ginras.ru/index/english/database_eng.html. Python source code implementing the cluster analysis of the earthquake catalog is available at https://github.com/SkurativskaKateryna/IJ_NLM_Earthquake_clusterization.git.
@@ Appendix A Estimation of DBSCAN parameters
The selection of the parameters $\varepsilon$ and Min_samplesMin\_samples for the DBSCAN algorithm is based on the analysis of the sorted $k$-distance graph, as discussed by [26] and [15], which uses the sorted distance matrix. We plot the $k$-distance graphs for $k=4$, $k=70$, and $k=160$ (Fig. 7). The appropriate value of $\varepsilon$ corresponds to the formation of the ”knee” on these graphs. Therefore, for our studies, we select $\varepsilon$ from the range $[0.01;0.2]$ and Min_samplesMin\_samples from the range $[10;300]$. Subsequent calculations indicate that the latter range can be narrowed to $[80,200]$, which is used throughout the paper. It is important to note that this procedure should still be considered an approximate method for parameter estimation, as noted by [16] and [54].
@@
Appendix B Fitting a mixture of normal distributions to earthquake depth data
To evaluate the PDFs plotted in Fig.6, we fit a mixture of normal distributions to the corresponding depth histograms for each cluster. To do this, we used the Python module sklearn.mixture containing the class GaussianMixture providing tools for working with Gaussian Mixture Models (GMMs). Particularly, we specify two component mixture of normals, i.e., $aN_{1}(\mu_{1},\sigma_{1})+bN_{2}(\mu_{2},\sigma_{2})$. The parameters $a$, $b$ ($a+b=1$), the expectations $\mu_{1,2}$, and the standard deviations $\sigma_{1,2}$ should be estimated. For instance, considering the total number of earthquakes, the depth histogram can be constructed (Fig.8), and corresponding parameters of the mixture of normals $a=0.7119$, $\mu_{1}=10.1612$, $\sigma_{1}=5.1950$, $b=0.2881$, $\mu_{2}=31.8807$, and $\sigma_{2}=14.99$ are calculated.
Figure 8: Fitting the mixture of normal modes to the earthquake depth histogram. The locations of modes (filled black dots) are 10.29 km and 31.63 km.
For the clusters obtained by DBSCAN, the calculated values of parameters are presented in Table 2. Corresponding PDF profiles are shown in Fig.6a. Considering the clusters obtained by HDBSCAN, we evaluated parameter values presented in Table 3 and plot the corresponding PDF profiles in Fig.6b.
Table 2: The parameters of the mixture of normal distributions (partition “db 192”) depicted in Fig.6a.
$a$ $\mu_{1}$ $\sigma_{1}$ $b$ $\mu_{2}$ $\sigma_{2}$
Cluster 1 0.5403 14.7111 5.141 0.4597 28.0154 5.4431
Cluster 2 0.5580 16.264 3.8686 0.442 8.6523 3.7392
Cluster 3 0.649 10.7328 3.596 0.351 5.2912 2.8054
Cluster 4 0.5357 5.5324 2.8294 0.4643 12.2318 3.7938
Cluster 5 0.7111 10.4547 2.2046 0.2889 5.9833 1.7462
Cluster 6 0.7061 8.961 2.3141 0.2939 13.9953 1.9496
Table 3: The parameters for the mixture of normal distributions (partition “hdb 15”) depicted in Fig.6b.
$a$ $\mu_{1}$ $\sigma_{1}$ $b$ $\mu_{2}$ $\sigma_{2}$
Cluster 1 0.6086 6.4897 3.1488 0.3914 13.6962 3.1542
Cluster 2 0.7563 10.2446 1.4319 0.2437 6.8023 1.3428
Cluster 3 0.5452 9.91 2.6428 0.4548 16.4508 2.7848
Cluster 4 0.5604 14.6678 4.3653 0.4396 25.7917 4.1998
Cluster 5 0.5915 10.1408 2.4734 0.4085 4.7844 2.2276
Cluster 6 0.8044 10.6700 3.1267 0.1956 3.2650 1.9019